OpenStudy (rhood.ra):

-3p 3 + 5p + (-2p 2) + (-4) - 12p + 5 - (-8p 3)

OpenStudy (adonine):

5p3 - 2p2 - 7p + 1

OpenStudy (welshfella):

@AdoNine PLeas don't give direct answers. Read the site's Code of Conduct.

OpenStudy (adonine):

ok..............

OpenStudy (gabbyalicorn):

Let's simplify: (−3p)(3)+5p+(−2p)(2)−4−12p+5−(−8p)(3) =−9p+5p+−4p+−4+−12p+5+24p Then, Combine Like Terms: =−9p+5p+−4p+−4+−12p+5+24p =(−9p+5p+−4p+−12p+24p)+(−4+5) =4p+1 You have your answer: =4p+1